Total fixed costs = $1,000,000 Unit Price = $5,515 Unit Variable Cost = $2,170 Find the breakeven volume. What happens to the breakeven volume if the unit price falls to $5,000 and unit variable cost rises to $2,500? Discuss your findings.
Answer #1

Answer

Let Quantity be Q. Total Revenue(TR) = Price*Quantity = 5515Q.

Total Cost(TC) = Fixed Cost + Total Variable cost = Fixed Cost + Average Variable cost*Quantity = 1,000,000 + 2170Q

Break even occurs when TR = TC

=> 5515Q = 1,000,000 + 2170Q

=> (5515 - 2170)Q = 1,000,000

=> Q = 298.95 ---------------------------Break even Volume.

Now Average variable cost increases from 2170 to 2500 and P decreases to 5000

Now New TC = 1,000,000 + 2500Q and TR = Total Revenue = 5000Q

Break even occurs when TR = TC

=> 5000Q = 1,000,000 + 2500Q

=> (5000 - 2500)Q = 1,000,000

=> Q = 400 --------------------------- New Break even Volume.

As Average Variable cost increases and Price decreases thus now more output should be produced in order to breakeven.